# More than Carbon: Cradle-to-Grave environmental impacts of GenAI training on the Nvidia A100 GPU

### v1

@@ Introduction
Utilization ratio Large-scale data centers typically divide their infrastructure into distinct partitions to accommodate different types of workloads - most commonly training, inference and other deep learning tasks [54]. We investigate a GPU cluster that is used exclusively for training purposes in the training partition of a data center. We estimate a GPU utilization_ratio of 100% for the training clusters. This assumption is justified by two key factors: first, training workloads are highly predictable compared to inference tasks, enabling effective resource planning; second, specialized AI training data centers deploy sophisticated workload schedulers specifically designed to minimize idle time and maximize utilization. While inference clusters may experience significant under-utilization due to unpredictable demand spikes, training partitions can sustain near-optimal utilization through proper workload management and exclusive GPU allocation methods used for large-scale foundational model training. This parameter will be subject to a SA.
Lifespan Based on extensive real-world data measurements in cluster computers [36, 14, 21] (detailed in SM), we adopt an average three-year lifespan for the A100 GPU in our study. This parameter will also be subject to a SA.
@@
The recycling and recovery of materials at the end-of-life (EoL) stage are considered using the formula and data developed by Ecosystem [44]. The allocation approach chosen is the cut-off method at the ‘no benefit’ substitution point. This approach is based on the principle that recycling or energy recovery of EoL materials does not provide benefits in terms of substituting virgin materials or primary energy sources. This implies that no environmental benefit is allocated at the EoL.
Despite the AI models being trained in two different countries with differing recycling strategies, the same EoL impact method is applied to both. This is due to data limitations on electronic waste streams and the proportion of materials collected, processed and recycled. Furthermore, the datasets under consideration exclusively encompass legal processes, while the environmental ramifications of illegal e-waste disposal are not included in the analysis [13].
@@ Results
The results are presented in two sections. First, we examine the Nvidia A100 SXM 40GB GPU as a physical product, generating intermediate results at the component and system level as illustrated in Fig. 1. Second, to assess the environmental impacts of AI-as-a-service, we scale these intermediate results according to two functional unit configurations: training BLOOM and GPT-4.
@@
In this section, intermediate impacts results are presented at the scale of the physical Nvidia A100 SXM 40GB GPU card throughout its lifecycle with an additional focus on the manufacturing stage. A usage context similar to BLOOM was considered with the following parameters : 3-year operational lifespan, assuming continuous operation at maximum capacity (100% utilization) with power consumption at 400W, connected to the French electricity grid in 2022. The impacts at the scale of the GPU card reveal variations in environmental burden distribution across the four life cycle stages, i.e., manufacturing, distribution, use and EoL, for 16 impact categories (see Fig. 3).
The use phase emerges as the dominant contributor across 11 out of 16 categories. Particularly for the impact category climate change, the use phase contributes to about 87% of total impact, despite the energy consumption being based on the French electricity mix that is predominantly based on nuclear energy. The use phase also contributes high shares to the categories resource use (fossil fuels) ($\geq$98%), ionizing radiation ($\geq$96%), EF-particulate matter ($\geq$95%), eutrophication (terrestrial) ($\geq$90%), and acidification ($\geq$79%) indicating the energy-intensive nature of the product’s operational stage. The elevated ionizing radiation impact can be attributed to the substantial contribution of nuclear energy to France’s electricity grid composition.
@@
The manufacturing stage represents the second most dominant life cycle stage, with substantial contributions to 5 out of 16 impact categories. Specifically, human toxicity, cancer ($\geq$99%), human toxicity, non-cancer ($\geq$56%), resource use, minerals and metals ($\geq$85%) and ecotoxicity, freshwater ($\geq$37%) . This reflects the material extraction and processing activities associated with product fabrication, particularly the use of metals and chemical substances in manufacturing processes. The substantial contribution to ozone depletion ($\geq$83%) can partially be explained by the release of HFC emissions and other ozone-depleting substances used in semiconductor manufacturing.
Even though the component transportation was modeled as air transport, distribution has relatively modest contributions across most categories, with its highest impact observed in eutrophication, marine ($\geq$2%) and eutrophication, terrestrial ($\geq$1%).
The EoL stage has its largest contribution to ecotoxicity, freshwater ($\geq$35%) and land use ($\geq$18%) [^3] We exclude the analysis of water use results for the EoL stage, as the background database is subject to error in this particular criterion.. However, these contributions should be interpreted with caution due to significant data limitations in modeling electronic waste streams. The environmental impacts are likely underestimated, as current LCA databases do not account for illegal and informal disposal practices that are prevalent in global e-waste management [13]. Additionally, the relatively modest contribution of the EoL stage compared to manufacturing and use phase reflects the relative dominance of the other life cycle stages rather than indicating low absolute environmental impacts from disposal.
The results demonstrate hotspots in the product’s life cycle, with the use phase representing the primary environmental burden for most impact categories, followed by manufacturing for material-intensive and toxicity impact categories and EoL for specific disposal-related categories.
Given that the manufacturing stage contributes substantially to the overall environmental impacts of a GPU card, following our analysis of impact distribution across all life cycle stages, we examine the manufacturing impacts per component for each impact category (see Fig. 4). The manufacturing stage encompasses the emissions associated with resource extractions, mineral processing, distribution and component manufacturing, collectively representing the cradle-to-gate stage. The cradle-to-gate impacts assessment of the GPU components reveals heterogeneity in environmental burden distribution across the five primary components: GPU chip, PCB, heatsink, PoP, casing and upstream distribution (see Fig.4).
@@
The GPU chip emerges as the dominant contributor across 10 out of 16 of impact categories and shows particularly pronounced contributions to climate change (81.8%), and resource use, fossils (80%), reflecting the energy intensive semiconductor fabrication processes required for GPU chip manufacturing compared to simpler ICs in the PCB. These high manufacturing impacts are further intensified by the A100’s advanced 7nm process technology, where smaller node size require increasingly energy and resource-intensive manufacturing processes per unit area. The transition to finer lithography processes necessitates more complex fabrication steps, higher precision manufacturing equipment, and greater energy consumption per cm2 of silicon area [40, 2], contributing to the relatively high environmental impact of the component. Regarding abiotic resource depletion of elements (ADPe) [^4] ADPe reflects the use of nonrenewable resources in relation to their scarcity., the main die does not show high impacts because gold is primarily associated with connections on the PCB or within specific types of IC packages (bonding) on the PCB rather than the main die itself. In contrast, the heatsink’s large copper content drives the overall resource depletion results for this category (80.1%), overshadowing the gold impact present in relatively small amounts within the PCB. Furthermore, the heatsink dominates human toxicity categories, contributing 91% to cancer-related impacts and 68.0% to non-cancer impacts. This can be attributed to the heatsink’s high copper content, comprising 98% of the component’s weight. As a heavy metal, which extraction and processing can further release toxic by-products [20], copper drives these toxicity impacts. Overall, the heatsink component shows impacts primarily in material-intensive categories, which is consistent with the extensive copper extraction and processing activities required for thermal management components. This component also contributes dominantly to eutrophication, freshwater (86%), ionising radiation, human health (90.7%), and land use (91%). These impacts likely stem from copper mining operations, which typically involve large-scale excavation (land use impacts), acid mine drainage contributing to surface water contamination [20] (accounting for eutrophication), and processing of ores containing naturally occurring radioactive materials [3] (ionising radiation impacts).
@@
The casing component exhibits the lowest overall environmental contribution, contributing less than 1% to most categories, indicating the relatively simple manufacturing processes required for GPU housing components. The manufacturing results for one A100 reveal clear environmental hotspots at the component level. Fabrication of the main die in semiconductor facilities represent the primary environmental burden for energy and climate-related impacts, while material processing activities for the heatsink and PCB dominate toxicity and resource depletion categories. This is consistent with other studies showing that ICs often represent a large share of manufacturing impacts [40].
To translate these intermediate product-level results into the impacts of one active GPU hour (our reference flow), we apply equation 4 enabling standardized comparison across different AI training scenarios and facilitating the subsequent scaling to our FU configurations to represent the environmental impacts of complete model training processes. The relative distribution of impact contributions per one GPU hour remains unchanged as shown in Figure 3; only the absolute environmental impacts per metric will be scaled down to reflect the impacts of one active GPU hour. This hourly metric then serves as the basis for estimating the final LCIA results according to equation 3.
@@
This section presents the LCIA results for our two FU configurations, which characterize the training of GPT-4 and BLOOM. This approach captures the environmental implications of AI-as-a-service considered at the GPU cluster scale, where computational resources are deployed temporarily for specific training objectives before being reallocated or decommissioned. Subsequently, the results are normalized following the planetary boundary framework. Finally, given that results show significant variability depending on operational assumptions, we conduct a sensitivity analysis to account for a range of usage contexts.
@@
The use phase dominates 11 out of 16 impact categories for training BLOOM, including climate change (86.9%), resource use, fossils (98.5%), and eutrophication, terrestrial (90.1%). The manufacturing stage accounts for substantial contributions across 5 of the 16 impact categories, more precisely resource use, minerals and metals (85.1%), and ozone depletion (83.2%). Further, the manufacturing stage contributes most to the toxicity related impact categories, human toxicity, cancer (99.6%), human toxicity, non-caner (56.6%), and ecotoxicity, freshwater (37.5%). The exact amounts can be found in the SM.
@@
This section presents the LCIA results of training BLOOM and GPT-4 normalized to the PBs per capita. This normalization translates complex LCA results into an intuitive framework that communicates environmental significance in the equivalence of the global ecological limits at the scale of one human inhabitant of the earth, making the results more accessible to policymakers and stakeholders outside the LCA community. To contextualize the LCA results within this framework, impacts are normalized to the PB-aware reference situation as provided by [45] for the year 2023 (world population of 8.01 billion [52]) corresponding to the year in which GPT-4 was trained [^5] Although BLOOM was trained in 2022, adjusting the normalization to the 2022 world population would alter the results by less than 1%. Consequently, the same normalization procedure is employed for both models..
@@
The average_power, lifespan and utilization_ratio of the cards are subject to significant variability depending on the context of use. To take this variability into account and represent a range of usage contexts, three sensitivity analyses are conducted on these three parameters through a min-max approach. The SA results are presented in Fig. 7 with the original scenario for each FU configuration serving as the reference point.
As established in equation 4 the average_power of the card exerts a direct and linear influence on its electricity consumption. This parameter depends upon several factors, including the intensity of utilization and the cooling of the GPUs. The reference scenario under consideration assumes an average power equivalent to the TDP of the GPU: 400W. The TDP of the PCIe version of the A100 can go up to 300W [33]. Conversely, even if a consumption of 500W is considerably higher than the standard TDP, this power can be attained within specific cooling designs as outlined in the Nvidia A100 Product documentation [33]. Thus, in the context of this SA, the parameter is varied between 300W and 500W.
As established in equation 1 the lifespan of the GPUs has an impact on the amortization of upstream (raw material acquisition and manufacturing) and downstream (EoL) impacts. The durability of the GPUs depends on a number of factors, including the intensity of utilization or the cards cooling. The reference scenario under consideration assumes a lifespan of 3 years. In the context of this SA, the parameter is varied to one year, two years, and four years.
The utilization_ratio of the card exerts an influence on both the electricity_consumption of the GPU (equation 4) and the amortization of its impacts over active GPU hours (equation 1). The reference scenario under consideration assumes a utilization ratio of 100% representing a maximalist approach. In the context of this SA, the parameter is varied between 85% and 50%. This parameter depends upon the mutualisation of the cards and the planning strategy that is employed.
Figure 7: Sensitivity analysis for GPT-4 and BLOOM. The reference scenario is defined with a 3 year lifespan, 400 W average power consumption and 100% utilization rate (author illustration).
@@
An increase, or a decrease, in power consumption is accompanied by a corresponding increase, or decrease, in the magnitude of impacts. This is particularly noteworthy for criteria that are strongly influenced by the use phase, such as climate change, particulate matter and the depletion of fossil resources, which vary between 75% and 125%. The deployment of energy efficiency measures is thus crucial in order to address the challenges associated with the impacts of the electrical grid (depending on the location of the data center).
For impact criteria influenced by the manufacturing phase, both lifespan and utilization_ratio strongly influence the results. The greatest variation is around 300% for human cancer toxicity for both GPT-4 and BLOOM, and nearly 300% for ionising radiation for GPT-4 with a one-year lifespan, reducing to 75% with a four-year lifespan. However, this evolution should not be interpreted as an actual impacts evolution at the level of the card, but rather at the cluster level. Increasing the lifespan of GPU cards reduces the need to replace them, resulting in fewer cards being replaced with new ones, which would have their own environmental impacts. Regarding the utilization_ratio, the increase in impacts should also be interpreted at the cluster level. Increasing the utilization_ratio reduces the number of cards needed to perform the same workload thanks to mutualization. In other words, the cluster will provide more GPU hours with the same quantity of cards by reducing the idle hours. This implies a reduction in the number of cards produced, each of which would have its own environmental impacts. Therefore, maximizing the utilization_ratio through mutualization and planning, as well as increasing the lifespan of the cards, seems crucial in order to reduce the impacts related to human and non-human toxicity, as well as mineral and metal resource depletion.
@@
However, higher utilization rates create sustained thermal stress that accelerate silicon degradation. This can manifest as decreased performance, increased latency, and a higher susceptibility to hardware errors over time [21]. In addition, high utilization - continuous near-maximum power draw - generates substantial heat loads that contribute to various failure mechanisms. These effects compound over time [36], potentially reducing the effective operational lifespan of the hardware and thereby increasing the amortized environmental impacts per unit of computational work performed. Even systems with modest aggregated utilization rates may experience prolonged periods of maximum power draw during training phases, subjecting the hardware to the same thermal stress as continuously high-utilization rates. This suggests that simply reducing allocation rates may not necessarily extend hardware lifespan or reduce life cycle environmental impacts. The reason is that AI training workloads are characterized by intensive computational bursts that push GPUs to their thermal limits regardless of overall utilization patters. During these training phases, GPUs operate at peak power consumption for extended periods, generating the same thermal stress and degradation mechanisms as would occur under continuous high utilization.
The interdependence between power consumption, thermal stress, and hardware lifespan suggests that optimization strategies should consider the full system dynamics rather than isolated metrics. However, because GPUs represent a significant capital investment, maximizing their allocation utilization is often a top priority for operators seeking efficient return on hardware expenditures. Yet, the economic optimization models used in practice do not account for the environmental costs of accelerated hardware replacement cycles. AI model developers and system architects can minimize the per-unit embodied environmental impact by ensuring maximum operational efficiency of system resources and extending the service life of AI infrastructure [53].
Discussion
This discussion contextualizes our findings within the broader landscape of AI environmental impacts research and examines the implications of our comprehensive LCA approach. We begin by validating our results against existing studies to assess consistency with scientific and industry-reported studies, highlighting the contribution of primary data collection. Finally, we address a critical gap in current AI sustainability discourse by examining environmental impacts beyond carbon emissions.
@@ Discussion
We discuss the consistency of our findings with previous estimates of GPU embodied impacts and AI training emissions through two studies: Nvidia’s HGX H100 carbon footprint report [34] and the life cycle impacts of BLOOM’s model training [26]. Given the absence of reproducible multi-criteria assessments of GPUs, these comparisons consider exclusively the climate change impact category.
@@
In 2025, Nvidia published a cradle-to-gate product carbon footprint (PCF) assessment of their HGX H100 system, reporting a total impact of 1,312 kg CO2eq. This assessment focuses exclusively on raw material extraction, mineral processing and component manufacturing for the climate change category only, enabling comparison with our intermediate LCIA results at the component level for the same impact category [34]. In the Nvidia report, the primary contributors to the carbon emissions are the materials and components, which represent 91% of the total footprint, while the assembly process accounts for an additional 8.6% and transportation contributes minimally at 0.4% to the overall footprint. For the materials and components category, the report provides a detailed breakdown by component type. High bandwidth memory emerges as the largest contributor (42%), followed by ICs (25%) and thermal components (18%). Nvidia also identifies electromechanical components (4%), common components (0.8%), PCBs (0.7%), mechanical components (0.6%), and interconnects (0.4%) as contributing to the overall PCF.
@@
The cradle-to-gate stage analysis from this study reveals a total carbon footprint of 127.6 kg CO2eq. for the single GPU, with the following distribution by component: GPU chip (main die and VRAM) contributes 82.3% to the total carbon footprint, representing the dominant emission source. To enhance comparability with the Nvidia study, we combine our main die, VRAM and PoP components (88.7%), which corresponds to their combined memory and IC categories (67%) [^7] While the PoP is not technically an IC, it is largely comparable to the components listed as ‘active components’ in the Nvidia report.. Our PCB analysis shows 6.4% contribution, which is 5.7% higher than Nvidia’s assessed PCB contribution to the overall carbon footprint. The substantial difference in system perimeter - GPU PCB in this study versus the HGX baseboard PCB in the Nvidia report - makes direct comparison uncertain. The thermal management components (heatsink) in our analysis contributes 3.8% of the carbon product footprint, approximately 15% lower than the assessment in the Nvidia report.
@@
Both studies demonstrate similar component hierarchy patterns, with semiconductor components (main die, VRAM, ICs) dominating the carbon footprint, followed by electronic components (PCB), and thermal management systems. The proportional distribution of emissions across component categories shows strong alignment between the two assessments, despite the architectural differences between the single A100 GPU and the HGX H100 baseboard system. Nevertheless, notable discrepancies emerge when accounting for memory capacity differences. The H100 incorporates 80GB of high bandwidth memory, which is double compared to our assessed GPU’s specifications, yet the observed emissions differential exceeds what memory scaling alone would predict. For one, this discrepancy might stem from methodological variations, particularly regarding data sources. Nvidia utilized ecoinvent 3.10 and Sphera’s LCA databases, whereas our assessment employed the Negaoctet database. Each database incorporates different modeling assumptions for semiconductor manufacturing processes. For the climate change category, there may be less variation between databases compared to other environmental indicators, although differences in system models and elementary flow can still contribute to discrepancies. Despite these methodological differences, both studies converge on identifying semiconductor components (specifically integrated circuits) as the primary carbon hotspot in GPU manufacturing, whether examined at the individual GPU scale or the server system level.

### v2

@@ Introduction
Utilization ratio Large-scale data centers typically divide their infrastructure into distinct partitions to accommodate different types of workloads - most commonly training, inference and other deep learning tasks [ye2024deep]. We investigate a GPU cluster that is used exclusively for training purposes in the training partition of a data center. We estimate a GPU utilization_ratio of 100% for the training clusters. This assumption is justified by two key factors: first, training workloads are highly predictable compared to inference tasks, enabling effective resource planning; second, specialized AI training data centers deploy sophisticated workload schedulers specifically designed to minimize idle time and maximize utilization. While inference clusters may experience significant under-utilization due to unpredictable demand spikes, training partitions can sustain near-optimal utilization through proper workload management and exclusive GPU allocation methods used for large-scale foundational model training. This parameter will be subject to a SA.
Lifespan Based on extensive real-world data measurements in cluster computers [Titan2020, meta2024llama3, karydopoulos2025heatimpact] (detailed in SM), we adopt an average three-year lifespan for the A100 GPU in our study. This parameter will also be subject to a SA.
@@
The recycling and recovery of materials at the end-of-life (EoL) stage are considered using the formula and data developed by Ecosystem [ecosystem]. The allocation approach chosen is the cut-off method at the ‘no benefit’ substitution point. This approach is based on the principle that recycling or energy recovery of EoL materials does not provide benefits in terms of substituting virgin materials or primary energy sources. This implies that no environmental benefit is allocated at the EoL.
Despite the AI models being trained in two different countries with differing recycling strategies, the same EoL impact method is applied to both. This is due to data limitations on electronic waste streams and the proportion of materials collected, processed and recycled. Furthermore, the datasets under consideration exclusively encompass legal processes, while the environmental ramifications of illegal e-waste disposal are not included in the analysis [Ficher2025].
@@ Results
The results are presented in two sections. First, we examine the Nvidia A100 SXM 40GB GPU as a physical product, generating intermediate results at the component and system level as illustrated in Fig. 1. Second, to assess the environmental impacts of AI-as-a-service, we scale these intermediate results according to two functional unit configurations: training BLOOM and GPT-4. Finally, we validate our results by comparing them with those of an industry report and a scientific study.
@@
We present intermediate impacts results at the scale of the individual Nvidia A100 SXM 40GB GPU, with detailed focus on the manufacturing stage. The assessment assumes operational parameters consistent with the BLOOM training infrastructure and context: 3-year operational lifespan, continuous operation at maximum capacity (100% utilization) with power consumption at 400W, connected to the French electricity grid (2022 carbon intensity). Environmental impacts vary at the scale of the GPU card across the four life cycle stages, i.e., manufacturing, distribution, use and end-of-life (EoL) and across 16 impact categories (see Fig. 3).
The use phase dominates 11 out of 16 categories. Particularly for the impact category climate change, the use phase contributes to about 87% of total impact, despite the energy consumption being based on the French electricity mix that is predominantly based on nuclear energy. The use phase similarly contributes high shares to the categories resource use (fossil fuels) ($\geq$98%), ionizing radiation ($\geq$96%), EF-particulate matter ($\geq$95%), eutrophication (terrestrial) ($\geq$90%), and acidification ($\geq$79%) reflecting the energy-intensive nature of GPU operations. The elevated ionizing radiation impact can be attributed to the substantial contribution of nuclear energy to France’s electricity grid composition.
@@
The manufacturing stage represents the second most dominant life cycle stage, with substantial contributions to 5 out of 16 impact categories. Specifically, human toxicity (cancer) ($\geq$99%), human toxicity (non-cancer) ($\geq$56%), resource use, minerals and metals ($\geq$85%) and ecotoxicity (freshwater) ($\geq$37%). This reflects the material extraction and processing activities associated with product fabrication, particularly the use of metals and chemical substances in manufacturing processes. The substantial contribution to ozone depletion ($\geq$83%) partially reflect the release of HFC emissions and other ozone-depleting substances during semiconductor manufacturing. Despite modeling the component transportation via air transport, the distribution contributes modestly across most categories, with its highest impact in marine eutrophication ($\geq$2%) and terrestrial eutrophication ($\geq$1%). The EoL stage contributes most substantially to freshwater ecotoxicity ($\geq$35%) and land use ($\geq$18%) [^4] We exclude the analysis of water use results for the EoL stage, as the background database is subject to error in this particular criterion.
However, these contributions should be interpreted with caution due to significant data limitations in modeling electronic waste streams. The environmental impacts are likely underestimated, as current LCA databases do not account for illegal and informal disposal practices that are prevalent in global e-waste management [Ficher2025]. Additionally, the relatively modest contribution of the EoL stage compared to manufacturing and use phase reflects the relative dominance of the other life cycle stages rather than indicating low absolute environmental impacts from disposal.
The results demonstrate hotspots in the product’s life cycle, with the use phase representing the primary environmental burden for most impact categories, followed by manufacturing for material-intensive and toxicity impact categories and EoL for specific disposal-related categories. Given manufacturing’s substantial contributes to the overall GPU environmental impacts, we further examine the manufacturing impacts per component across all impact category (see Fig. 4). The manufacturing stage encompasses the emissions associated with resource extractions, mineral processing, distribution and component manufacturing, collectively representing the cradle-to-gate stage. The component-level cradle-to-gate impacts assessment reveals heterogeneous environmental burden distribution across five primary elements: GPU chip, printed circuit board (PCB), heatsink, power-on-package (PoP), casing, and upstream distribution (see Fig.4).
@@
The GPU chip emerges as the dominant contributor across 10 out of 16 of impact categories and shows particularly pronounced contributions to climate change (81.8%), and resource use, fossils (80%), reflecting the energy intensive semiconductor fabrication processes required for GPU chip manufacturing compared to simpler ICs in the PCB. These high manufacturing impacts are further intensified by the A100’s advanced 7nm process technology, where smaller node size require increasingly energy and resource-intensive manufacturing processes per unit area. The transition to finer lithography processes necessitates more complex fabrication steps, higher precision manufacturing equipment, and greater energy consumption per cm2 of silicon area [Pirson2023_ICproduction, Boakes2023-jm], contributing to the relatively high environmental impact of the component. Regarding abiotic resource depletion of elements (ADPe) [^5] ADPe reflects the use of nonrenewable resources in relation to their scarcity., the main die does not show high impacts because gold is primarily associated with connections on the PCB or within specific types of IC packages (bonding) on the PCB rather than the main die itself. In contrast, the heatsink’s large copper content drives the overall resource depletion results for this category (80.1%), overshadowing the gold impact present in relatively small amounts within the PCB. Furthermore, the heatsink dominates human toxicity categories, contributing 91% to cancer-related impacts and 68.0% to non-cancer impacts. This can be attributed to the heatsink’s high copper content, comprising 98% of the component’s weight. As a heavy metal, which extraction and processing can further release toxic by-products [izydorczyk2021copper], copper drives these toxicity impacts. Overall, the heatsink component shows impacts primarily in material-intensive categories, which is consistent with the extensive copper extraction and processing activities required for thermal management components. This component also contributes dominantly to eutrophication, freshwater (86%), ionising radiation, human health (90.7%), and land use (91%). These impacts likely stem from copper mining operations, which typically involve large-scale excavation (land use impacts), acid mine drainage contributing to surface water contamination [izydorczyk2021copper] (accounting for eutrophication), and processing of ores containing naturally occurring radioactive materials [chau2008naturalradiat_copper] (ionising radiation impacts).
@@
The casing component exhibits the lowest overall environmental contribution, contributing less than 1% to most categories, indicating the relatively simple manufacturing processes required for GPU housing components. The manufacturing results for one A100 reveal clear environmental hotspots at the component level. Fabrication of the main die in semiconductor facilities represent the primary environmental burden for energy and climate-related impacts, while material processing activities for the heatsink and PCB dominate toxicity and resource depletion categories. This is consistent with other studies showing that ICs often represent a large share of manufacturing impacts [Pirson2023_ICproduction].
To translate these intermediate product-level results into the impacts of one active GPU hour (our reference flow), we apply equation 4 enabling standardized comparison across different AI training scenarios and facilitating the subsequent scaling to our FU configurations to represent the environmental impacts of complete model training processes. The relative distribution of impact contributions per one GPU hour remains unchanged as shown in Figure 3; only the absolute environmental impacts per metric will be scaled down to reflect the impacts of one active GPU hour. This hourly metric then serves as the basis for estimating the final life cycle impact assessment (LCIA) results according to equation 3.
@@
We present the LCIA results for two FU configurations, which characterize the training of GPT-4 and BLOOM. This approach captures the environmental implications of AI-as-a-service considered at the GPU cluster scale, where computational resources are deployed temporarily for specific training objectives before being reallocated or decommissioned. Subsequently, the results are normalized following the planetary boundary (PB) framework to contextualize impact magnitudes relative to Earth system thresholds. Given substantial variability in environmental outcomes across operational parameters, we conduct sensitivity analyses spanning realistic ranges of hardware utilization, lifespan, and power consumption to capture context-dependent uncertainty in impact estimates.
@@
The use phase dominates 11 out of 16 impact categories for training BLOOM, including climate change (86.9%), resource use, fossils (98.5%), and eutrophication, terrestrial (90.1%). The manufacturing stage accounts for substantial contributions across 5 of the 16 impact categories, more precisely resource use, minerals and metals (85.1%), and ozone depletion (83.2%). Further, the manufacturing stage contributes most to the toxicity related impact categories, human toxicity, cancer (99.6%), human toxicity, non-caner (56.6%), and freshwater ecotoxicity (37.5%). The exact amounts can be found in the SM.
@@
Normalizing the LCIA results of training BLOOM and GPT-4 to the planetary boundaries (PBs) per capita translates the complex LCA results into an intuitive framework that communicates environmental significance in the equivalence of the global ecological limits at the scale of one human inhabitant of the earth. This makes the results more accessible to policymakers and stakeholders outside the LCA community. To contextualize the LCA results within this framework, impacts are normalized to the PB-aware reference situation as provided by [sala2018_PB_LC_development] for the year 2023 (world population of 8.01 billion [WorldBank2023Pop]) corresponding to the year in which GPT-4 was trained [^6] Although BLOOM was trained in 2022, adjusting the normalization to the 2022 world population would alter the results by less than 1%. Consequently, the same normalization procedure is employed for both models..
@@
The GPU average power, lifespan and utilization ratio are subject to significant variability depending on the deployment context, directly influencing environmental impact assessments. To capture this variability, we conducted sensitivity analyses (SA) on these three parameters using a min-max approach spanning realistic operational ranges. The average power consumption exerts a direct linear influence on electricity consumption (Equation 4) and depends among others on utilization intensity and cooling system design. Our reference scenario assumes an average power equivalent to the GPU’s TDP of 400W. The PCIe version of the NVIDIA A100 operates at 300 W TDP, while specialized cooling configurations can sustain power draw up to 500 [NvidiaA100Product]. We therefore varied this parameter between 300 W and 500 W to encompass standard and high-performance operational conditions.
GPU lifespan affects the amortization of upstream impacts (raw material acquisition and manufacturing) and downstream end-of-life burdens across total GPU hours (Equation 1). Hardware durability depends on several factors, including utilization intensity and thermal management effectiveness. Our reference scenario assumes a 3-year lifespan, with sensitivity analysis spanning one to four years to reflect variability in replacement cycles across different operational contexts.
@@
The utilization ratio influences both electricity consumption (Equation 4) and impact amortization over active GPU hours (equation 1). Our reference scenario assumes 100% utilization, representing a maximalist approach, with sensitivity analysis ranging from 50% to 85% to reflect realistic operational variability. An increase, or a decrease, in power consumption is accompanied by a corresponding increase, or decrease, in the magnitude of impacts. This is particularly noteworthy for criteria that are strongly influenced by the use phase, such as climate change, particulate matter and the depletion of fossil resources, which vary between 75% and 125%. These findings underscore the importance of energy efficiency interventions for mitigating grid-related environmental impacts, with effectiveness contingent on data center location and electricity supply characteristics.
For manufacturing phase dominated impact categories, both lifespan and utilization ratio exert substantial influence on results. Human cancer toxicity exhibits the greatest variability, reaching approximately 300% for both GPT-4 and BLOOM, while ionizing radiation for GPT-4 varies from nearly 300% at one-year lifespan to 75% at four years. However, these variations reflect cluster-level impacts rather than individual card-level dynamics. Extended lifespan reduces replacement frequency, thereby decreasing the number of newly manufactured cards and their associated environmental burdens. Similarly increased utilization ratio, reduces at the cluster level the total number of cards required to deliver equivalent computational capacity through more efficient resource sharing, with clusters providing more GPU hours from the same hardware inventory by minimizing idle time. Both mechanisms ultimately reduce aggregate hardware production and its embodied impacts. These findings demonstrate that maximizing utilization through workload optimization and hardware sharing, coupled with extended replacement cycles, represents a crucial strategy for mitigating human, non-human and ecological toxicity impacts alongside mineral and metal resource depletion associated with AI infrastructure.
Figure 7: Sensitivity analysis for GPT-4 and BLOOM. The reference scenario is defined with a 3-year lifespan, 400 W average power consumption and 100% utilization rate (author illustration).
@@
However, higher utilization rates create sustained thermal stress that accelerate silicon degradation. This can manifest as decreased performance, increased latency, and a higher susceptibility to hardware errors over time [karydopoulos2025heatimpact]. In addition, high utilization - continuous near-maximum power draw - generates substantial heat loads that contribute to various failure mechanisms. These effects compound over time [Titan2020], potentially reducing the effective operational lifespan of the hardware and thereby increasing the amortized environmental impacts per unit of computational work performed. Even systems with modest aggregated utilization rates may experience prolonged periods of maximum power draw during training phases, subjecting the hardware to the same thermal stress as continuously high-utilization rates. This suggests that simply reducing allocation rates may not necessarily extend hardware lifespan or reduce life cycle environmental impacts. The reason is that AI training workloads are characterized by intensive computational bursts that push GPUs to their thermal limits regardless of overall utilization patters. During these training phases, GPUs operate at peak power consumption for extended periods, generating the same thermal stress and degradation mechanisms as would occur under continuous high utilization.
The interdependence between power consumption, thermal stress, and hardware lifespan suggests that optimization strategies should consider the full system dynamics rather than isolated metrics. However, because GPUs represent a significant capital investment, maximizing their allocation utilization is often a top priority for operators seeking efficient return on hardware expenditures. Yet, the economic optimization models used in practice do not account for the environmental costs of accelerated hardware replacement cycles. AI model developers and system architects can minimize the per-unit embodied environmental impact by ensuring maximum operational efficiency of system resources and extending the service life of AI infrastructure [wu2022sustainable].
@@
We discuss the consistency of our findings with previous estimates of GPU embodied impacts and AI training emissions through two studies: Nvidia’s HGX H100 carbon footprint report [NvidiaH100PCF] and the life cycle impacts of BLOOM’s model training [luccioni2023]. Given the absence of reproducible multi-criteria assessments of GPUs, these comparisons consider exclusively the climate change impact category.
@@
Nvidia’s 2025 cradle-to-gate product carbon footprint (PCF) for the HGX H100 system reports 1,312 kg CO2eq [NvidiaH100PCF]. The assessment focuses exclusively on raw material extraction, mineral processing and component manufacturing for the climate change category only, enabling comparison with our intermediate LCIA results at the component level for the same impact category. The primary contributors to the carbon emissions are the materials and components, which represent 91% of the total footprint, while the assembly process accounts for an additional 8.6% and transportation contributes minimally at 0.4% to the overall footprint. For the materials and components category, the report provides a detailed breakdown by component type. High bandwidth memory emerges as the largest contributor (42%), followed by ICs (25%) and thermal components (18%). Nvidia also identifies electromechanical components (4%), common components (0.8%), PCBs (0.7%), mechanical components (0.6%), and interconnects (0.4%) as contributing to the overall PCF.
@@
Our cradle-to-gate analysis of the Nvidia A100 GPU estimates 127.6 kg CO2eq per GPU, dominated by the GPU chip and VRAM (82.3%), followed by the PCB (6.4%) and heatsink (3.8%). When grouped analogously, semiconductor components (main die, VRAM, PoP) represent 88.7% of the total, aligning closely with Nvidia’s 67% for memory and ICs [^8] While the PoP is not technically an IC, it is largely comparable to the components listed as ‘active components’ in the Nvidia report.. Our PCB analysis shows 6.4% contribution, which is 5.7% higher than Nvidia’s assessed PCB contribution to the overall carbon footprint. The substantial difference in system perimeter - GPU PCB in this study versus the HGX baseboard PCB in the Nvidia report - makes direct comparison uncertain. The thermal management components (heatsink) in our analysis contributes 3.8% of the carbon product footprint, approximately 15% lower than the assessment in the Nvidia report.
Despite architectural and methodological differences, most notably system scope (single GPU vs. multi-GPU system), lithography (7 nm vs. 5 nm), and database choice (Negaoctet vs. ecoinvent/Sphera), the proportional distribution of emissions across component categories shows strong alignment between the two assessments and both studies identify semiconductor components (specifically integrated circuits) as the dominant carbon hotspot in GPU production.
@@
Nevertheless, notable discrepancies emerge when accounting for memory capacity differences. The H100 incorporates 80GB of high bandwidth memory, which is double compared to our assessed GPU’s specifications, yet the observed emissions differential exceeds what memory scaling alone would predict. For one, this discrepancy might stem from methodological variations, particularly regarding data sources. Nvidia utilized ecoinvent 3.10 and Sphera’s LCA databases, whereas our assessment employed the Negaoctet database. Each database incorporates different modeling assumptions for semiconductor manufacturing processes. For the climate change category, there may be less variation between databases compared to other environmental indicators, although differences in system models and elementary flow can still contribute to discrepancies. Despite these methodological differences, both studies converge on identifying semiconductor components (specifically integrated circuits) as the primary carbon hotspot in GPU manufacturing, whether examined at the individual GPU scale or the server system level.